# LEP-QNN: Loan Eligibility Prediction Using Quantum Neural Networks

### v1

@@ Abstract
LEP-QNN: Loan Eligibility Prediction Using Quantum Neural Networks
@@ I Introduction
Machine Learning (ML) technologies have conducted in a new era in predictive analytics, furnishing tools with the capability to learn from data and make informed predictions. These models find applications across a spectrum of financial operations, from credit scoring and fraud detection to the nuanced task of predicting loan eligibility [1, 2, 3, 4]. However, despite their considerable successes, these conventional ML paradigms are tethered to the intrinsic computational confines of classical computing, especially when confronted with the complexity of large-scale financial datasets [5, 6]. This computational impasse, exacerbated by the burgeoning volume and diversity of financial data, calls for a leap into computational frameworks that stride beyond the limitations of conventional computing methods.
Quantum Machine Learning (QML) stands at this frontier as a groundbreaking integration of Quantum Computing (QC) with ML algorithms [7, 8]. This innovative integration uses the principles of quantum mechanics to enhance ML models’ processing power and efficiency [9, 10, 11], opening up possibilities across various domains, from materials science to healthcare [12, 13, 14, 15]. Notably, its application within the finance sector has demonstrated profound impacts [16, 17, 18], underscoring the versatility and efficacy of QML in addressing complex, data-intensive challenges.
@@
  • A dropout mechanism is seamlessly integrated into the QNN-LEP framework, aimed at reducing overfitting and boosting the framework’s ability to generalize across diverse data sets.
@@
Our LEP-QNN framework embodies a significant advancement toward achieving more accurate, efficient, and inclusive financial services. By forging a nexus between QC and financial analytics, our study makes a notable contribution to the theoretical landscape of QML and underscores its practical implications in revolutionizing the financial industry.
@@ II-A QC and QML in Finance
The practical application of QML for solving real-world financial problems was further highlighted by resources offering guidance on exploiting quantum algorithms within a Python environment [27]. This approach underscored the advantages of QC in financial analysis and modeling, advocating for an integrated approach to harnessing the full power of quantum and classical algorithms. In credit scoring and classification, the efficiency of feature selection was notably enhanced by applying Quadratic Unconstrained Binary Optimization (QUBO) models on quantum annealers [28]. The exploration of fairness in QNNs introduced a balanced approach to deploying fair and accurate QNNs in finance [29]. This research highlighted the importance of incorporating fairness into the accuracy-driven domain of QC, proposing comprehensive solutions that balance both objectives. Finally, developing a QC method for calculating the value at risk for portfolios demonstrated the method’s effectiveness [30], showcasing the practical applicability of quantum technologies in accurately managing large asset portfolios.
@@ III-A QNN Architecture
The QNN within our LEP-QNN framework is shown in Fig. 3, showcasing the methodical arrangement of quantum operations that form the core of our computational model. The circuit is initiated through angle encoding, which effectively transposes each feature of the input data $x_{i}$ into quantum states:
@@
Following the angle encoding, the QNN utilizes an ansatz—a predetermined yet flexible structure of quantum gates—each parameterized by a set of variables $\theta$. These gates are responsible for manipulating the qubits to model complex nonlinear relationships within the data:
@@
where $U_{d}(\theta_{d})$ corresponds to the unitary operation at each depth $d$, where $d=5$ to outline the model’s multilayered complexity, with $\theta_{d}$ representing the set of parameters at that specific layer.
@@ IV-A Experimental Setup
Our experiments are conducted using a high-performance computing environment equipped with an Intel(R) Xeon(R) Platinum 8259CL CPU @ 2.50GHz, complemented by access to QC cloud services via the respective software tools, ensuring the rigorous and robust evaluation of our LEP-QNN framework.
@@ IV-D Comparison with Existing Works
The comparison of existing approaches, as detailed in Table III, highlights the groundbreaking accuracy achieved by our LEP-QNN framework. With an accuracy of 95%, our framework significantly outperforms conventional models using the same dataset, marking a substantial advance in predictive analytics for loan eligibility. The nearest competitor, an ensemble method comprising the best three classical algorithms, lags by 7.74%—a clear indication of the quantum model’s superior capability in managing complex data. This comparison attests to the novelty of our approach, given that prior works have not explored QML models for this dataset and loan prediction in general. The superior performance of LEP-QNN paves the way for future explorations into quantum models in the financial sector and beyond.
@@
Reference Approaches Accuracy (%) [37] Ensemble with Best Three 87.26 [38] Extra Trees 86.2 [4] Decision Tree with AdaBoost 84 [39] Naïve Bayesian 80.4 Our approach LEP-QNN 95
@@ V Conclusion
Concluding this study on QML applied to financial predictive analytics, our LEP-QNN framework demonstrates the significant impact of QC in enhancing the accuracy of financial models. With an impressive 98% At the core of our work is the development of a specialized QNN architecture. The inclusion of a unique dropout mechanism within this architecture is a critical advancement in overcoming overfitting, enhancing the model’s generalization across diverse financial scenarios. This feature underscores the practical value of quantum-enhanced models in real-world applications. Our comprehensive evaluation of different optimizers has shed light on optimal strategies to enhance QNN performance. This aspect of our research guided us to achieve optimum model efficiency and highlighted the delicate balance between quantum algorithms and performance enhancement. Moreover, examining the model’s endurance against various quantum noise models offers insights into its stability and reliability within the unpredictable nature of QC environments. As we wrap up, our LEP-QNN framework marks a milestone for financial analytics, showcasing QC as the key to innovation and precision in the field. This work opens doors to future explorations in QML, inviting further research into quantum technologies’ potential to address complex analytical challenges across different sectors. Through this study, we lay the groundwork for future advancements in QC applications, encouraging ongoing exploration of quantum technologies in analytics and beyond.

### v2

@@ I Introduction
Machine Learning (ML) technologies have conducted in a new era in predictive analytics, furnishing tools with the capability to learn from data and make informed predictions. These models find applications across a spectrum of financial operations, from credit scoring and fraud detection to the nuanced task of predicting loan eligibility [1, 2, 3, 4]. However, despite their considerable successes, these conventional ML paradigms are limited by the intrinsic computational confines of classical computing, especially when confronted with the complexity of large-scale financial datasets [5, 6]. This computational impasse, exacerbated by the burgeoning volume and diversity of financial data, calls for a leap into computational frameworks that stride beyond the limitations of conventional computing methods.
Quantum Machine Learning (QML) stands at this frontier as a groundbreaking integration of Quantum Computing (QC) with ML algorithms [7, 8]. This innovative integration uses the principles of quantum mechanics to enhance ML models’ processing power and efficiency [9, 10, 11], opening up possibilities across various domains, from materials science to healthcare [12, 13, 14, 15, 16, 17]. Notably, its application within the finance sector has demonstrated profound impacts [18, 19, 20], underscoring the versatility and efficacy of QML in addressing complex, data-intensive challenges.
@@
  • A dropout mechanism is seamlessly integrated into the LEP-QNN framework, aimed at reducing overfitting and boosting the framework’s ability to generalize across diverse datasets.
@@
Our LEP-QNN framework embodies a significant advancement toward achieving more accurate, efficient, and inclusive financial services. By forging a nexus between QC and financial analytics, our study makes a notable contribution to the theoretical landscape of QML and underscores its practical implications in revolutionizing the financial industry. This paper is organized as follows. In Sec. II, we review background and related work on QC in finance. In Sec. III, we introduce the LEP-QNN framework and its components. In Sec. IV, we present results and discussions on optimizer performance and noise model analysis. Finally, in Sec. V, we conclude with a summary of findings and future directions.
@@ II-A QC and QML in Finance
The practical application of QML for solving real-world financial problems was further highlighted by resources offering guidance on exploiting quantum algorithms within a Python environment [29]. This approach underscored the advantages of QC in financial analysis and modeling, advocating for an integrated approach to harnessing the full power of quantum and classical algorithms. In credit scoring and classification, the efficiency of feature selection was notably enhanced by applying Quadratic Unconstrained Binary Optimization (QUBO) models on quantum annealers [30].
The exploration of fairness in QNNs introduced a balanced approach to deploying fair and accurate QNNs in finance [31]. This research highlighted the importance of incorporating fairness into the accuracy-driven domain of QC, proposing comprehensive solutions that balance both objectives. Finally, developing a QC method for calculating the value at risk for portfolios demonstrated the method’s effectiveness [32], showcasing the practical applicability of quantum technologies in accurately managing large asset portfolios.
@@ III-A QNN Architecture
The QNN within our LEP-QNN framework is presented in Fig. 3, showcasing the methodical arrangement of quantum operations that form the core of our computational model. The circuit is initiated through angle encoding, which effectively transposes each feature of the input data $x_{i}$ into quantum states:
@@
Following the angle encoding, the QNN utilizes an ansatz predetermined yet flexible structure of quantum gates—each parameterized by a set of variables $\theta$. These gates are responsible for manipulating the qubits to model complex nonlinear relationships within the data:
@@
where $U_{d}(\theta_{d})$ corresponds to the unitary operation at each depth $d$, where $d=5$ to outline the model’s multilayered complexity, with $\theta_{d}$ representing the set of parameters at that specific layer. In designing the ansatz, we review several configurations inspired by existing literature. We implement a pre-testing phase—an algorithmic evaluation that systematically examines different ansatz candidates. In this phase, the following steps are performed:
  • Configuration Sampling: A set of ansatz candidates is generated, each differing in gate composition and layering. These candidates include various combinations of rotation gates, controlled operations, and entangling layers, designed to capture non-linearities in a diverse manner.
  • Algorithmic Evaluation: Each candidate is subjected to an initial training cycle using a portion of the loan eligibility dataset. We use performance metrics such as prediction accuracy, convergence speed, and loss reduction as benchmarks. Cross-validation and iterative testing are employed to ensure that the evaluation is robust and that performance remains consistent across multiple data splits.
  • Systematic Comparison: The performance of each ansatz candidate is compared using a scoring metric that combines accuracy in predicting loan eligibility with the model’s capability to represent the essential features of the data.
Through this systematic comparison, we identify the fixed ansatz structure as the most effective option for our framework. This choice is based on its superior performance in achieving higher prediction accuracy and its alignment with our methodological objectives—namely, representing the essential features of the data pertinent to loan eligibility. Moreover, the chosen ansatz offers flexibility and can be refined further.
@@ IV-A Experimental Setup
Figure 5: The experimental setup in our evaluation.
Our experiments are conducted using a high-performance computing environment equipped with an Intel(R) Xeon(R) Platinum 8259CL CPU @ 2.50GHz, complemented by access to QC cloud services via the respective software tools, ensuring the rigorous and robust evaluation of our LEP-QNN framework (see Fig. 5).
@@ IV-D Comparison with Existing Works
The comparison of existing approaches, as detailed in Table III, highlights the groundbreaking accuracy achieved by our LEP-QNN framework. With an accuracy of 98%, our framework significantly outperforms conventional models using the same dataset, marking a substantial advance in predictive analytics for loan eligibility. The nearest competitor, an ensemble method comprising the best three classical algorithms, lags by 7.74%—a clear indication of the quantum model’s superior capability in managing complex data. This comparison attests to the novelty of our approach, given that prior works have not explored QML models for this dataset and loan prediction in general. The superior performance of LEP-QNN paves the way for future explorations into quantum models in the financial sector and beyond.
@@
Reference Approaches Accuracy (%) [39] Ensemble with Best Three 87.26 [40] Extra Trees 86.2 [4] Decision Tree with AdaBoost 84 [41] Naïve Bayesian 80.4 Our approach LEP-QNN 98
@@ V Conclusion
Concluding this study on QML applied to financial predictive analytics, our LEP-QNN framework demonstrates the significant impact of QC in enhancing the accuracy of financial models. With an impressive 98% accuracy rate in loan eligibility predictions, LEP-QNN highlights the advanced capabilities of QML in processing complex financial data and establishes new benchmarks for applying quantum technologies in data analysis. At the core of our work is the development of a specialized QNN architecture. The inclusion of a unique dropout mechanism within this architecture is a critical advancement in overcoming overfitting, enhancing the model’s generalization across diverse financial scenarios. This feature underscores the practical value of quantum-enhanced models in real-world applications. Our comprehensive evaluation of different optimizers has shed light on optimal strategies to enhance QNN performance. This aspect of our research guided us to achieve optimum model efficiency and highlighted the delicate balance between quantum algorithms and performance enhancement. Moreover, examining the model’s endurance against various quantum noise models offers insights into its stability and reliability within the unpredictable nature of QC environments. As we wrap up, our LEP-QNN framework marks a milestone for financial analytics, showcasing QC as the key to innovation and precision in the field. This work opens doors to future explorations in QML, inviting further research into quantum technologies’ potential to address complex analytical challenges across different sectors. Through this study, we lay the groundwork for future advancements in QC applications, encouraging ongoing exploration of quantum technologies in analytics and beyond.